The sum of two numbers is eight. If seven times the smaller number is negative five times the larger number, find the numbers.
Answer by Herdaysoji(25)

first number=x
second number=y
x+y=8, but7y=-5x, 5x+7y=0
simultaneous equation
x+y=8
5x+7y=0
multiply with the coefficients of x(which are 5 and 1) and introduce negative to second equation
5*(x+y=8)=5x+5y=40.......(equ 1)
1*(5x+7y=0)=-(5x+7y=0)...... (equ 2)
.'.5x cancels -5x
so, 5y=40
-7y=0
I.e-2y=40 , (divide both side by -2),y=-20
substitute y into equation 1
x+y=8, x+(-20)=8
x-20=8, x=8+20, I.e x=28
so x=28, y=-20